# Scaling Laws for Function Diversity and Specialization Across Socioeconomic and Biological Complex Systems

### v2

@@ 2.3 Model predictions and categorization of systems by diversification and specialization parameters
While the specialization parameter, $\gamma$, exhibits common patterns across systems in terms of its mean value, the diversification parameter, $\theta$, differs significantly across systems. It is notably higher for cities than for federal agencies and cells. A higher $\theta$ value suggests the existence of a highly abundant function suppresses new function creation. This is in line with the observation that large cities tend to become hubs for certain industries [60], such as San Francisco Bay Area for information technology and the Greater Boston area for biotech. A value of $\theta$ around zero implies that an increased range of existing functions, regardless of the size of the functions, contributes towards suppression of new functions. The differences in $\theta$ parameters can help explain the differences in scaling behavior observed. When $\theta=1$, given that $\sum_{i=1}^{D}k_{i}=N$, we can simplify Eq. 2 to derive $p=p_{0}/N$. Since $p=dD/dN$, we arrive at the dynamical equation $dD/dN=aD_{\text{max}}/N$. The solution of this equation takes the form of $D\sim\log(N)$, which recovers the logarithmic scaling. Similarly, in the case of $\theta=0$, we have $\sum_{i=1}^{D}k_{i}^{0}=D$. This leads to $dD/dN\sim 1/D$, which implies the power-law scaling $D\sim N^{1/2}$. For bacteria, Figure 1(A) shows that diversification happens faster than in cities, but slower than in federal agencies, and correspondingly the mean calibrated value of $\theta$ is inbetween 0 and 1.
@@
Each class of systems occupies distinct regions of parameter space. First, the values of $\theta$ found for cities imply that diversity growth is most impeded when large functions are abundant. Second, for federal agencies, smaller $\theta$ values imply that functions more equitably impede diversity growth. Finally, for bacterial cells, an intermediate $\theta$ value implies an in-between level of diversity growth still impeded by large functions. Cities express the largest value of $\theta$ possible (approximately 1) that simultaneously allows for the maximization of concentration inzq few functions and universality of the rank-frequency distributions without leading to gelation behavior that occurs for $\gamma>1$ [37]. In this sense, gelation provides a natural bound to function growth since it does not allow for the robustness or diversity necessary for organizations to survive complex environments.
3 Discussion
@@
We identify common patterns in the diversity and abundance of functions across a broad spectrum of biological and social complex systems. In systems such as bacteria, federal agencies, companies, and universities, function diversity can be approximated by a power-law relationship with size, characterized by sublinear exponents ranging from 0.35 to 0.57, depending on the system. This is consistent with Heaps’ Law. However, the function diversity of cities follows a different pattern, scaling logarithmically with size.
We propose a general model for the growth of function diversity and abundance that accounts for both power-law and logarithmic scaling behaviors. The model is characterized by two key parameters: a diversification parameter, which governs the expansion of function range, and a specialization parameter, which grows function concentration. We estimate both parameters from the data for bacterial cells, federal agencies, and cities. Notably, we observe consistent sublinear preferential across all three systems, with generally small variations in the specialization parameter. However, the diversification parameter varies significantly across systems, suggesting different drivers for function range expansion. In our model, the probability of new function addition is, by definition, a decreasing function of the size of the organism. This choice is motivated by research indicating that not only is this the case in biological systems [61, 62], but more recently as a principle of ecology in the decreasing relationship between speciation rates and animal size (see [63, Figure 3]).
@@
There are several limitations to the proposed model. First, the choice of a modified Yule-Simon type process to model diversity growth is likely not unique. Compared to previous models [51], our model contains an extra parameter, which effectively allows modulation between the probability of adding new functions compared to the strength of preferential attachment. This is necessary since it does not seem possible to unify the dynamics of diversity for federal agencies, cells, and cities using a model with fewer than two parameters. The classic Yule-Simon model, which assumes a constant rate of new function addition, leads to a linear scaling of diversity with system size which is not observed in the systems we consider (which are either sublinear power-law or logarithmic). Frameworks under other narratives, such as urn models or sample-space reducing mechanisms, could potentially also be adapted with an extra parameter for the same purpose. While our model considers the addition of individuals, individuals also leave organizations and cities, just as genes can naturally be removed from genomes through evolution. It could be useful for future research to consider the removal of individuals in complex systems and how it would affect diversification and specialization. Additionally, our model assumes a direct proportionality between system size and age, though this may not always hold. However, the relationship is generally expected to be monotonically increasing on average, as described by Cope’s rule [68]. Our proposed mechanism is an effective mechanism that can describe the emergence of structures from a unifying, albeit simplified, framework. This provides a way to produce the diverse observed empirical pattern across systems in a minimal way, to observe whether coordination problems across biology and society utilize similar solutions.
Cities differ significantly from bacterial cells and federal agencies across various empirical measures. Unlike the power-law scaling observed in bacteria and federal agencies, cities exhibit logarithmic scaling in diversity. Additionally, cities display greater universality in abundance distributions. This divergence may be attributed to the fact that cities are structurally distinct from the other two systems. Federal agencies and bacterial cells function as integrated units with well-defined boundaries, akin to organisms, and the goal for adapting to its environment is on the level of each organism. In contrast, cities resemble ecosystems, lacking a singular objective, with diverse entities operating together, each primarily based on self-interest rather than collective adaptation. This fundamental difference in structure and goal underpins the unique scaling behaviors observed in cities, distinguishing them from the more organism-like dynamics of bacterial cells and federal agencies.
@@
Our paper makes several contributions to the literature. First, in contrast to the prevalent trend of studying diversification and specialization within specific system types, our work presents an integrated analysis across a diverse array of systems, encompassing both biological (bacteria) and social systems (companies, government agencies, universities, cities). This approach identifies commonalities across many biological and social systems, while also identifying notable exceptions. It provides testable hypotheses to explore patterns in other systems. Second, while previous models explained power-law or logarithmic scaling relationships using distinct formulations and assumptions, we develop a general model that explains both types of scaling within one unifying framework. Third, by estimating parameters in the framework of this mathematical model, we can quantify common characteristics across complex systems and evaluate how systems differ from each other.
Acknowledgements
The authors would like to acknowledge the support of the National Science Foundation Grant Award Number EF–2133863. We would like to thank the UNM Center for Advanced Research Computing, supported in part by the National Science Foundation, for providing the research computing resources used in this work, in particular, Matthew Fricke. H.Y. acknowledges the Emergent Political Economies grant from the Omidyar Network, the NRF Global Humanities and Social Sciences Convergence Research Program (2024S1A5C3A02042671) and the support from the Institute of Management Research at Seoul National University. J.H. acknowledges the support of a Lou Schuyler grant from the Santa Fe Institute.
@@ Acknowledgements
The code necessary to replicate the results of the calibrations, and simulate the model, is available at https://github.com/jamesholehouse/Calibrations-for-Function-Diversity.

### v3

@@ Results
At the same time, patterns emerge across systems along both $\theta$ and $\gamma$, revealing shared mechanisms that operate both within and across system classes. For instance, bacteria, federal agencies, and cities exhibit parallel trends in the specialization parameter $\gamma$, which captures how the abundance of existing functions shapes subsequent growth. All three systems consistently show positive $\gamma$ values, but predominantly sublinear (below 1), indicating a general tendency toward diminishing returns in specialization. Nevertheless, the variation in $\gamma$ is wider for federal agencies and bacteria than for cities, suggesting that urban systems may exhibit a more uniform—and possibly universal—specialization pattern.
To test whether the parameter patterns observed for cities are robust to temporal and contextual variation, we repeated the calibration across multiple years for the four largest urban areas (Figure 5). The results show no systematic temporal trends, despite population of these cities having changed significantly over this time period—indicating that both $\theta$ and $\gamma$ remain stable over time. We also perform robustness tests and demonstrate the stability of these parameters to data noise, initial conditions, and alternative assumptions (see SI).
@@
While the specialization parameter, $\gamma$, exhibits common patterns across systems in terms of its mean value, the diversification parameter, $\theta$, differs significantly across systems. It is notably higher for cities than for federal agencies and cells. A higher $\theta$ value suggests the existence of a highly abundant function suppresses new function creation. This is in line with the observation that large cities tend to become hubs for certain industries [7], such as San Francisco Bay Area for information technology and the Greater Boston area for biotech. A value of $\theta$ around zero implies that an increased range of existing functions, regardless of the size of the functions, contributes towards suppression of new functions. The differences in $\theta$ parameters can help explain the differences in scaling behavior observed. When $\theta=1$, given that $\sum_{i=1}^{D}k_{i}=N$, we can simplify Eq. 1 to derive $p=p_{0}/N$. Since $p=dD/dN$, we arrive at the dynamical equation $dD/dN=aD_{\text{max}}/N$. The solution of this equation takes the form of $D\sim\log(N)$, which recovers the logarithmic scaling. Similarly, in the case of $\theta=0$, we have $\sum_{i=1}^{D}k_{i}^{0}=D$. This leads to $dD/dN\sim 1/D$, which implies the power-law scaling $D\sim N^{1/2}$. For bacteria, Figure 1(A) shows that diversification happens faster than in cities, but slower than in federal agencies, and correspondingly the mean calibrated value of $\theta$ is between 0 and 1.
@@
Each class of systems occupies distinct regions of parameter space. First, the values of $\theta$ found for cities imply that diversity growth is most impeded when large functions are abundant. Second, for federal agencies, smaller $\theta$ values imply that functions more equitably impede diversity growth. Finally, for bacterial cells, an intermediate $\theta$ value implies an in-between level of diversity growth still impeded by large functions. Cities express the largest value of $\theta$ (approximately 1) that simultaneously allows for the maximization of concentration in a few functions while maintaining consistent rank-frequency distributions without leading to gelation behavior that occurs for $\gamma>1$ [43]. In this sense, gelation provides a natural bound to function growth since it does not allow for the robustness or diversity necessary for organizations to survive complex environments.
By enabling cross-domain comparisons, we identify important commonalities and differences across complex systems. The barrier to introducing new functions varies across system types. Yet, once a function is added, it tends to grow following a common sublinear preferential attachment dynamics. The variations may stem from differences in system goals and structures. Federal agencies and bacterial cells function as cohesive organisms with well-defined boundaries, where adaptation occurs at the level of the whole entity. In federal agencies, the introduction of new occupations is typically decided in a centralized manner. Cities, on the other hand, lack unified objectives and resemble ecosystems: decentralized, and adaptation occurs at the level of the individuals. The introduction of new occupational categories is decided in a decentralized fashion. For organism-like systems, it is beneficial to develop a full set of functional capabilities to effectively adapt to environmental challenges. Thus, even when some functions dominate, there remains a drive to cover all essential bases. In cities, however, large functions may inhibit the emergence of others due to competitive or self-reinforcing dynamics. Once functions are introduced into these systems, the growth of function abundance follows a shared pattern across all systems: sublinear preferential attachment. This suggests that while the drivers of function creation differ based on system architecture and goals, the growth of functions once introduced follows a remarkably universal process.
Figure 5: Calibrated parameters ($\theta$ and $\gamma$) for the four largest metropolitan areas in the longitudinal dataset, by year. Parameter values show no systematic temporal trends.
Discussion
@@
We identify common patterns in the diversity and abundance of functions across a broad spectrum of biological and social complex systems. In systems such as bacteria, federal agencies, companies, and universities, function diversity can be approximated by a power-law relationship with size, characterized by sublinear exponents ranging from 0.35 to 0.57, depending on the system. This is consistent with Heaps’ Law. However, cities differ significantly from these systems and exhibit logarithmic scaling in function diversity and greater commonality in abundance distributions. This empirical anomaly may stem from cities’ fundamentally different structures and goals.
To reconcile these empirical patterns, we propose a general model for the dynamics of function diversity and abundance. The model is characterized by two fundamental parameters: a diversification parameter, which governs the expansion of the function range, and a specialization parameter, which grows the function concentration. We estimate both parameters from the data for bacterial cells, federal agencies, and cities. Notably, we observe consistent sublinear preferential attachment across all three systems, with generally small variations in the specialization parameter. However, the diversification parameter varies significantly across systems, suggesting different drivers for function range expansion. We emphasize that this universality lies at the level of generative mechanism rather than identical empirical outcomes. For simplicity, the model that we propose is fully Markovian, although future extensions to non-Markovian models with memory [25, 65] would be a beneficial addition to the literature, especially in the context of memory of organizational routines. The tight clustering of cities in parameter space implies a high degree of structural regularity in how urban systems specialize and diversify. This consistency supports the view that urban economies may be shaped by shared underlying mechanisms, as proposed in the urban scaling literature [15]. Extending this analysis to cities across different geographical and governance contexts would provide a useful test of the model’s generality.
@@
A key to this cross-domain comparison is the definition of diversity. Our use of the term in this paper specifically refers to the number of unique functions in the sub-components within a system. This definition differs from entropy-based measures stemming from information theory, which have also been used as a diversity index in the biological and social systems literature. Entropy-based measures combine two properties of distributions—richness and evenness—into a single index. While valuable in many contexts, entropy-based measures have the limitation of possessing mathematical properties that do not align with theoretical or intuitive understandings of diversity in applied domains [39]. In contrast, our diversity metric solely measures richness, the system’s range of unique capabilities. We explicitly separate the dimension of evenness, which we refer to as specialization, into a distinct second axis.
Our analysis relies on classification systems that delineate what constitutes a distinct function in each domain, which have limitations. For example, a protein’s role can sometimes vary depending on the cellular context across species, most retain consistent biochemical functions that permit meaningful functional classification. Social systems exhibit similar variations within function categories: occupations are defined by shared core tasks even as their implementation varies across organizations. The “chief executive” occupation category, for example, contains leaders of both start-ups and multinational corporations, whose daily activities differ, yet they are grouped together because they share core tasks of strategic oversight and coordination. The protein–occupation analogy thus reflects commonality in functional organization across diverse contexts rather than a literal one-to-one correspondence. Admittedly, classification systems, especially those in social systems, are shaped by administrative needs. However, they are updated regularly to reflect evolving labor and knowledge structures, and are widely used in real-world economic and policy decisions. We use the most granular level of these taxonomies to approximate a system’s repertoire of task specializations. Though no classification is perfect, the persistence of scaling patterns across domains that use different classification schemes suggests that the underlying dynamics of functional diversification and specialization reflect deeper regularities in system organization despite limitations in the classification systems.
Although proteins can exhibit context-dependent roles across species, most retain consistent biochemical functions that permit meaningful functional classification. Social systems show a comparable pattern: occupations are defined by shared core tasks even as their implementation varies across organizations. A start-up founder and a corporate CEO, for example, differ in daily activities yet share core functions of coordination and strategic control, and are thus categorized under the same SOC category. The protein–occupation analogy thus reflects a common structure of functional organization across diverse contexts rather than a literal one-to-one correspondence
Our work provides a parsimonious framework for unifying observed patterns across domains, but naturally abstracts away many details. This opens several directions for extension. While our model accounts for cross-sectional differences in scaling behavior across system types, future work could extend this framework to incorporate temporal dynamics, considering the evolution of the diversification and specialization parameters ($\theta$, $\gamma$) over time at different stages in system development. Our proposed framework represents an effective process, but does not explicitly capture the causal dynamics through which new functions or components emerge. Future work could extend this framework by explicitly modeling the structural dependencies among functions through hierarchical trees or network-based mechanisms. These include hierarchical specialization (depth), recombination across existing types (horizontal linkage), and the introduction of entirely novel types (breadth). Another promising extension involves accounting for function obsolescence. Our model focuses on the addition of individuals and functions, but individuals also leave organizations and cities, just as genes can naturally be removed from genomes through evolution. Entire functions can also become obsolete, such as the obsolescence of switchboard operators who manually connected telephone calls. Although the mechanisms of agent removal and function obsolescence are still areas of active research [45, 67], it could be useful for future research to consider the removal of individuals and functions in complex systems and how it would affect diversification and specialization. Lastly, our model treats the introduction of new functions as a single process, without distinguishing between the creation of globally novel functions and the adoption of ones that exist in other entities. In both biological and social systems, new functions arise within an entity and spread through diffusion—for example, via horizontal gene transfer in biological systems [3] and imitation in social systems [21]—reflecting a shared innovation–diffusion dynamic. Because the adoption of functions that exist in other organisms is likely far more common than the creation of globally new ones, it may dominate the observed empirical patterns. Future work could explicitly separate these processes to assess their relative contributions.
@@
Our work makes several contributions to the literature. First, in contrast to the prevalent trend of studying diversification and specialization within specific system types, our work presents an integrated analysis across a diverse array of systems, encompassing both biological (bacteria) and social systems (companies, government agencies, universities, cities). This approach identifies commonalities across many biological and social systems, while also identifying notable exceptions. It provides testable hypotheses to explore patterns in other systems. Second, while previous models explained power-law or logarithmic scaling relationships using distinct formulations and assumptions, we develop a general model that explains both types of scaling within one unifying framework. Third, by estimating parameters in the framework of this mathematical model, we can quantify common characteristics across complex systems and evaluate how systems differ from each other.
@@
The data and code necessary to replicate the results of this study are available at
https://github.com/jamesholehouse/Calibrations-for-Function-Diversity.
Author contributions
V.C.Y. and J.H. contributed equally to this work. Data gathering and empirical analysis: V.C.Y., J.I.A.; Model development, simulation, and calibration: J.H., S.R.; Writing – Initial Draft: V.C.Y, J.H.; Writing – Review & Editing: All authors; Conceptualization: All authors.
Acknowledgments
The authors would like to acknowledge the support of the National Science Foundation Grant Award Number EF–2133863. We would like to thank the UNM Center for Advanced Research Computing, in particular, Matthew Fricke, supported in part by the National Science Foundation, for providing the research computing resources used in this work. H.Y. acknowledges the Emergent Political Economies grant from the Omidyar Network, the NRF Global Humanities and Social Sciences Convergence Research Program (2024S1A5C3A02042671) and the support from the Institute of Management Research at Seoul National University. J.H. acknowledges the support of a Lou Schuyler grant from the Santa Fe Institute.